# Minimizing PLM-Based Few-Shot Intent Detectors

### v1

@@ Abstract
Haode Zhang1  Xiao-Ming Wu1thanks:   Corresponding author.  Albert Y.S. Lam2 Department of Computing, The Hong Kong Polytechnic University, Hong Kong S.A.R.1 Fano Labs, Hong Kong S.A.R.2 haode.zhang@connect.polyu.hk, xiao-ming.wu@polyu.edu.hk, albert@fano.ai
@@ 1 Introduction
Intent detection refers to the task of classifying users’ utterances according to the underlying intents. The task is critical since the result determines the following procedures such as dialogue state tracking, and thus the final output of the dialogue system. Pre-trained language models (PLMs) make it easier to train an intent detector with only a few labeled data, but these models are difficult to deploy in resource-constrained scenarios such as mobile devices due to the gigantic model sizes.
@@ Methods.
We use the following LLMs to generate data. GPTJ-6B Wang and Komatsuzaki (2021) with 6 billion parameters. OPT-30B Zhang et al. (2022) with 30 billion parameters. GPT4 OpenAI (2023), one of the most popular LLMs  [^2] We use the GPT4 variation of text-davinci-003.. We also experiment with another two student model architectures: convolutional neural networks (CNN) Chen (2015) and BiLSTM Graves and Schmidhuber (2005).

### v2

@@ Abstract
Haode Zhang1  Albert Y.S. Lam2  Xiao-Ming Wu1thanks:   Corresponding author. Department of Computing, The Hong Kong Polytechnic University, Hong Kong S.A.R.1 Fano Labs, Hong Kong S.A.R.2 haode.zhang@connect.polyu.hk, xiao-ming.wu@polyu.edu.hk, albert@fano.ai
@@ 1 Introduction
Intent detection refers to the task of classifying users’ utterances according to the underlying intents. The task is critical since the result determines the following procedures such as dialogue state tracking, and thus the final output of the dialogue system. Pre-trained language models (PLMs) make it easier to train an intent detector with only a few labeled data Zhang et al. (2021, 2022a), but these models are difficult to deploy in resource-constrained scenarios such as mobile devices due to the gigantic model sizes.
@@ Methods.
We use the following LLMs to generate data. GPTJ-6B Wang and Komatsuzaki (2021) with 6 billion parameters. OPT-30B Zhang et al. (2022b) with 30 billion parameters. GPT4 OpenAI (2023), one of the most popular LLMs  [^2] We use the GPT4 variation of text-davinci-003.. We also experiment with another two student model architectures: convolutional neural networks (CNN) Chen (2015) and BiLSTM Graves and Schmidhuber (2005).